OpenStudy (anonymous):

Joe is performing a dilation. One of his points is Y'(-4, 2), and another point is Y(-2, 1). What scale factor is he using? 2 -1/2 1/2 1

OpenStudy (anonymous):

It is not 1/2

hero (hero):

I forget which is the image and which is the pre-image. I think its 2

hero (hero):

If it's not 1/2, then it has to be 2.

OpenStudy (anonymous):

The pre-image is the original image the image is the new image

Directrix (directrix):

I'm betting on 1:2 assuming Y' is the post image and y the pre-image.

hero (hero):

then that means Y is the original image and Y' is the dilated image. Directrix, you have it backwards.

Directrix (directrix):

-2/x = 1/2 x = -4 where x is the x-coordinate of the post-image or image

OpenStudy (anonymous):

I have alreaady tried 1/2 it is not right

Directrix (directrix):

Okay. I concur with Hero's 2 but I can't guarantee anything for myself.

Directrix (directrix):

It cannot be 1. If 1/2 is out, -1/2 is also out. That leaves 2.

OpenStudy (anonymous):

2 was right

hero (hero):

Where's my medal?

Directrix (directrix):

Your medal?

OpenStudy (anonymous):

There is your medal

hero (hero):

lol

Directrix (directrix):

@ Hero --> Check this. Based on my stumbling here, I think we know this answer. http://openstudy.com/study#/updates/4f4e4b7ee4b019d0ebae284a

OpenStudy (anonymous):

Ill give you one 2 Directrix

Directrix (directrix):

Thank you.